40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog, Page 39 of 56  Not logged in ELOG logo
ID Date Author Type Category Subject
  888   Mon Aug 2 00:46:13 2010 DmassLaserPSLH1 NPRO Measured

The second order RC low pass filter does this according to LISO -

since you can't treat a second order system of RCs like this as though it were modular and just cascade them, it seemed prudent to include a liso plot:

 

Vin___4.7k________4.7k________Vout

                        |  1n              |  1n

                      gnd               gnd

Attachment 1: Picture_2.png
Picture_2.png
  887   Sun Aug 1 23:16:45 2010 DmassLab InfrastructureGeneralWHERE IS THE WINCAMD?

I was trying to understand why the mode matching through the PMC was so bad (I got 30 mW transmitted with 120 mW input).

  • I looked at the mode on the PMC REFL camera, and it did not look very guassian.
  • I changed the laser current from 2 A to 2.2 A (what Rick said they ran it at at LHO) and saw no change in the mode
  • I tried to take a beamscan...but could not find the WINCAMD and its computer (again.) I checked the usual suspects (PSL/TCS) and didn't see it.

So. Who knows where the wincamD is?

added: pics of PMC REFL before and after locking, in order

Totally possible that my mode matching is just that crappy? Maybe.

I did take beam scans of the mode going into the PMC just before aligning to it for the final time. I can do so again to get real profiles and characterize the H1NPRO

Attachment 1: IMG_0248.png
IMG_0248.png
Attachment 2: IMG_0250.jpg
IMG_0250.jpg
  886   Sun Aug 1 18:37:07 2010 DmassLaserPSLH1 NPRO Measured

I found a bad 6dB attenuator which I had in rotation, it seems to either be mislabeled, or have a broken connection. I labeled it NFG?

Audio Locking:

I am using 270 kHz at about 5 mV into the PMC to generate audio sidebands. I chose 270 because it has around a modulo pi phase, which means I don't have to build a phase shifter, and it seemed to have an acceptable FM / AM ratio.

 

  885   Fri Jul 30 19:11:10 2010 AlastairLaserGYRORubidium

The measurements in this plot were made using the 13dB LO mixer.  We increased the gain on the SR560 going into the daq to 1000 and AC coupled it to get above the daq noise floor.

Measurement 1 was made when the phase of the clocks was not quite exactly pi/2.  The issue here is that we can be sensitive to amplitude fluctuations as well as phase.  The output voltage from the mixer was at -200mV.  We then swapped the bnc cable lengths to get as close as possible to a pi/2 phase difference.  In the end we got to 20mV (the calibration here is 1v/pi rads and at 0v the signals have pi/2 phase).

Rana noticed that according to the clock's spec they are pretty sensitive to temperature fluctuations.  The third measurement in these plots was with some 'thermal insulation' on the clocks.  We didn't leave them long enough to reach any sort of equilibrium temperature, but it should have blocked out the external fluctuations a little.  Koji and Jenne then stole the thermal insulation from our experiment - thus putting personal comfort above scientific discovery (though Koji at least did elog what time they did it).

I changed the mixer in the phase measurement for a 7dB local oscillator one at around 7pm.  Data from before this time should still be calibrated using the 1V / pi rads value.  At this point I also had to change the gain on the SR560 to 100 to stop it railing.

When I changed the mixer this evening, I also tried unplugging the 1pps signal that is locking the clocks.  There was no noticeable drift, and I think we could easily make a measurement by leaving them synced and then unplugging them just before we measure.

Rana also suggested that we try making making a measurement with a phase difference of 0 rad, to make an amplitude noise measurement.

RA: I changed most of the 'pi's' above to 'pi/2'. The phase noise measurement is best done with a 90 deg phase shift between the signals. Either 0 or pi gives an optimum for amplitude noise measurements. Alastair changed the mixer from a Level 13 ZP-3MH mixer to a Level 7 XXXX mixer because the 10 MHz output of the Rb clocks has an amplitude of ~450 mVrms into 50 Ohms (that's pretty close to +7 dBm). In the previous setup, the mixer was working in a somewhat non-ideal mode so we can't really trust the results. We will have to take back Jenne and Koji's jackets and retry the measurement.

Attachment 1: rubidium_measurement.pdf
rubidium_measurement.pdf
  884   Fri Jul 30 16:48:11 2010 DmassLaserPSLH1 NPRO Measured

I measured the capacitance of the BLK-89S+ from minicircuits. They include neither a cut off frequency, nor the capacitance in the datasheet. Just a picture of a capacitor.

220 nF is the answer I got.

Thorlabs PD - 50 Ohm output -----220 nF----->mixer input (50 Ohms to ground?) seems to imply that the DC block has a transfer function of S/2/(S+1/2RC), with a frequency of 7 kHz. Everything gets through.

  883   Fri Jul 30 15:36:22 2010 DmassComputingDAQElog changes to the framebuilder

I will soon be taking data with the framebuilder. Please elog restarts / etc...If  dataviewer is frozen when I come in I want to know if someone restarted some process, or if something about the frontend is messed up

  882   Fri Jul 30 13:38:56 2010 DmassLaserDoublingMode Matching for Ovens

I redid the mode matching into the ovens to account for the dispersion of the material affecting the Rayleigh range (like this)

I used Zr = L_crystal / 2 / n

Attachment 1: OvenMM.pdf
OvenMM.pdf
  881   Fri Jul 30 01:56:07 2010 DmassLab InfrastructurePurchasesBought some lenses

I bought some lenses (of the AR.18 variety) to replace the missing ones from the newport lens kit. People should continue to do this. They will likely arrive while im at LLO, so someone should put them in the kit when they get here.

  880   Fri Jul 30 00:48:11 2010 DmassLaserGYROSeismometer Measurements Continued

Quote:

I've added SR560s to all three channels of the seismometer now.  They're DC coupled with a gain of 100, low pass filter at 10k and I'm running them on battery to avoid grounding the seismometer outputs.  The manual says that the battery should last 20hours so we might need to use a different solution to get more data (I'm assuming we don't want to use the mini circuits bnc transformers for this since it's not RF?).

We are now well above the DAQ noise.  Here is a plot showing 10 averages of the vertical, N/S and E/W data plotted calibrated in volts (GPS 964463447).  We will now take seismic data for 24hours.  Setup completed at 11.40am July29.

 

The bottom 560 was making a bunch of clicking noises and had flashing red lights on the overload and battery indicators. I have turned it off.

  879   Thu Jul 29 20:05:05 2010 AlastairLaserGYRORubidium clocks

 I couldn't get data from dataviewer earlier, so I came over to the 40m to look at it here. Unfortunately the system here is down at the moment and Alberto is rebooting it just now.

I went in to take a look at the beat signal on a scope and the good news is that the clocks have now synced.  I'll get back out here tomorrow some time to take some data.  They are not at the zero-crossing point, but are sitting about 200mV off (500mV is pi radians of phase).  I'll try to adjust this with some different lengths of BNC tomorrow.

Till I get proper read-out, here is a picture of the scope screen.

Attachment 1: rubidium.JPG
rubidium.JPG
  878   Thu Jul 29 19:19:42 2010 AlastairLaserGYROCW beam alignment

 This afternoon I tried getting the clockwise beam (AOM beam) into the cavity.  We didn't have much success the last time, mainly because it was so difficult aligning the beam from waaaaaayy back before the faraday.  So this time I tried a different approach.  The plan this time was to align the CW beam to the cavity, and then to put the faraday back in place after the alignment.

I took out the faraday altogether, and the mirrors after the faraday.  I then aligned the incoming CW beam along a line of holes on the bench and checked the 4" height along the length.  This step was just so that we had a well defined beam path to put the faraday back in place later.

After some fiddling with the mirrors that go after the faraday I got the CW beam aligned to the cavity and resonating.  I made sure it was overlapping the CCW beam in the cavity, and checked that the reflected CW beam was overlapping along the input path.  Then I put the faraday back in place and moved it around till I got the beam through it and the CCW reflected beam back out of it.  I put the 1/2 wave plate back in, checked the polarization and then setup the mirror, lens and PD for the reflection locking of the CCW beam.  I got the cavity locked in the CCW direction again after changing the power back up to 5mW in that direction (somewhere between 5 and 10mW seems to give a nice error signal.  It may lock below this, but we were using 5mW before so I'm continuing with that just now).

With the CCW locked on the 00 mode I then swept through the frequency on the AOM till I got the 00 for the CW direction.  Then I played with the alignment a little till this looked good (I've not tried maximizing the transmitted power in this direction yet like we did for the CCW direction, so that's on the 'to do' list).  Then I got the PD for the reflection locking of this direction aligned and put the signal into the second PDH box.  I checked that we were getting an error signal by sweeping the cavity and we are.  Then I hooked the PDH box up to the frequency modulation input of the VCO for the AOM.  

The CW direction was already close to being resonant on the 00 mode so it was difficult to tell if this was really locking.  As a sanity check I t-ed the output of the PDH box to a scope and then changed the center frequency on the VCO for the AOM.  As I changed the frequency you could see the output of the 2nd PDH box tracking this and the CW direction of the cavity stayed locked to the 00 mode. 

Attachment 1: cavity.jpg
cavity.jpg
  877   Thu Jul 29 13:20:13 2010 AlastairComputingGeneralDataviewer

I'm not sure if this is a known issue, but I can't get dataviewer to plot channels from the 40m when I am at a workstation in the ATF.  I can get data using DTT, and I can get dataviewer to connect (using nodus.ligo.caltech.edu and port 8088). I can see the channels listed but when I try either realtime or playback it won't show the data.

  876   Thu Jul 29 12:09:56 2010 AlastairLaserGYROSeismometer Measurements Continued

I've added SR560s to all three channels of the seismometer now.  They're DC coupled with a gain of 100, low pass filter at 10k and I'm running them on battery to avoid grounding the seismometer outputs.  The manual says that the battery should last 20hours so we might need to use a different solution to get more data (I'm assuming we don't want to use the mini circuits bnc transformers for this since it's not RF?).

We are now well above the DAQ noise.  Here is a plot showing 10 averages of the vertical, N/S and E/W data plotted calibrated in volts (GPS 964463447).  We will now take seismic data for 24hours.  Setup completed at 11.40am July29.

Attachment 1: 10_07_29_seismic01.pdf
10_07_29_seismic01.pdf
  875   Wed Jul 28 21:18:44 2010 AlastairLaserGYRORubidium clock

 The new rubidium clock arrived today from SRS.  It is a FS725 and the calibration sheet, manual and control program on floppy are in the filing cabinet in the ATF.

At the moment we are doing some testing of the clock, beating it against a second rubidium clock at the 40m.  The setup is that we have the 10MHz output feeding into a mixer (it is a 13dBm mixer at the moment but we should change this).  Each output is transformer coupled to avoid ground loops.  The output of the mixer goes through a low pass filter (10MHz I believe) and into an SR560 with a gain of 1 and DC coupling.  This signal is going into the channel mc_drum1 of the ADC.

The clocks are drifting relative to each other with several minutes period.  We need to make the measurement of phase noise close to the zero crossing of the beat so that we are sensitive to phase and not amplitude.  We did a few things to try to get the clocks locked in frequency.

First we tried locking the clocks using the freq adjust input on the back of the new clock.  We took the mixer output (giving ~500mV for pi radians phase difference) and used an SR560 to feedback.  It wasn't clear whether this was doing anything or not.  The frequencies did get closer over ~30mins, but this may just have been because the new clock is still aging (and also warming up).  After this time it appeared not to be getting much closer.

Next we removed the calibration sticker to get at the manual frequency adjust underneath it, and gave 3 clockwise turns (0.0025Hz per turn).  There was no immediate (or noticeable) change to the beat frequency.

We have now left them with the newer clock synced using the 1pps output from the old clock.  The sync light is on, but the frequencies were not getting any closer when I left them.  We're taking data and we'll look at it again tomorrow to see if they are getting closer.  There is some long time constant used for this locking loop (~2hours) that will make this take quite a long time.   We can shorten this time constant using the software and RS232 input if we need to.

 

 

 

 

 

  874   Wed Jul 28 19:51:10 2010 DmassLaserPSLH1 NPRO Measured

Quote:

Here is what I think is a transfer function of the frequency response of the H1NPRO PZT. Ideally, I would use a resonance with modulo 180 degrees
 of phase shift, and a relatively flat phase / frequency response

RA:  No units on the Y-axis???

 DYM - Not yet, since I am getting the signal for the transfer function by mixing down the beats between the various sidebands, the calibration wasn't immediately obvious to me (though it may well be simple).

Here is a diagram of the measurement.

N.B. there is an attenuator right after the first power splitter on its way to the NPRO PZT, I am not giving it 7 dBm of juice

Attachment 1: IMG_0241.png
IMG_0241.png
  873   Tue Jul 27 19:04:35 2010 DmassLaserPSLH1 NPRO Measured

Here is what I think is a transfer function of the frequency response of the H1NPRO PZT. Ideally, I would use a resonance with modulo 180 degrees
 of phase shift, and a relatively flat phase / frequency response

RA:  No units on the Y-axis???

Attachment 1: NproPZTFMBode.pdf
NproPZTFMBode.pdf
  872   Tue Jul 27 14:46:42 2010 JennaLaserGYROSeismometer Measurements

We have the Gurlap seismometer working, and it took data overnight (mostly) successfully. The North-South channel was not turned on, but as of this morning all three channels (North-South, East-West, and Vertical) are up and running. I've attached a plot of some data for the E-W and Vert channels comparing the noise at 3am and 11am. We calibrated the data by first converting from counts to m/s by multiplying by 40V/2^16*(1/802V/(m/s)). The 802V/(m/s) calibration comes from the Guralp spec sheet (which I've attached as a pdf). The calibration for all three channels is different, but the difference is less than 1%, so I've just chosen the calibration value for East-West and used it for all channels.

The channel names are

Vertical-- C2:ATF-GENERIC_DOF6_OUT_DAQ

North-South-- C2:ATF-GENERIC_DOF7_OUT_DAQ

East-West-- C2:ATF-GENERIC_DOF8_OUT_DAQ

After we have 24 hours of data, we'll make plots for more times on all three channels.

Attachment 1: seismic727.pdf
seismic727.pdf
Attachment 2: Guralp_spec_sheet
  871   Tue Jul 27 00:10:20 2010 DmassLaserPSLPMC Calibrated

I could sweep a couple FSR using a function generator into the PMC. A few numbers:

  • PMC FSR - 713 MHz (From design doc)
  • Voltage at drive box for 1 FSR = 3.5 +/- 0.3 V
    • Error bars generated via nonlinearity: (I ran through three FSR on one sweep, with one of the FSR around 0V input, the two FSR were 3.2 and 3.8 Volts, repsectively).
  •  
  • Voltage at PZT with 0V into driver = 142 V
  • 10 Vpp triangle wave at drive box = 184 Vpp at PZT drive (measured using a 10 Vpp triangle wave out of a function generator, watching the PZT voltage with a Fluke)
  • Calibration at input to drive box - 202 +/- 20MHz / Volt
  • Calibration at PZT - 11.1 MHz +/- 1 MHz/ Volt

Spectra shall be calibrated henceforth, or at the very least, calibratable.

Added: Calibrated PMC control signal in Hz/rtHz- As this should be totally coherent with err, I should just be able to divide by the open loop transfer function between ERR and CTRL...(I guess my Hz -> Volts bit of the OLTF on the error signal needs to be measured to get this no matter what...so I would end up with a Volts ->Hz for my error signal anyways...)

  • Foton isn't letting me add filters for some reason. It didn't earlier today, and Aidan had to change permissions (via chmod) to some process (Which?), then I could. I left and came back, and I could no longer write to Foton again.
  • Note that I am feeding back on the temperature of the NPRO.
  • WHOOPS: For correct spectrum I need to apply 2 poles @ 1Hz, 2 zeros at 10 Hz, 1 pole at 72 Hz, and a unity DC gain...
Attachment 1: PMCCtrl.pdf
PMCCtrl.pdf
  870   Fri Jul 23 00:45:46 2010 DmassLaserPSLH1 NPRO Measured

I locked using the PZT pole at 70 Hz with an SR560 - I am using a feed through with the front end to generate an integrator with a pole at 70 Hz and add it to this signal with the other 560 input channel

I ran into some problems seeing anything in DTT. Frank can see all his channels (C3 PSL), but we can't seem to get the C2 ATF) channels. Will investigate further.

The transfer function:

  • Dewhitening in Driver Box:
    • 2 zeros at 10 Hz
    • 2 poles at 1 Hz
  • PMC PZT Pole from driver box / PZT capacitance
    • 72 Hz
  • 2f filtration
    • 2 poles at 10 kHz
  869   Thu Jul 22 21:27:23 2010 ZachLaserGYROGYRO update

Not that I see how this could have changed, but did you do the standard error signal checks? If you switched a cable here or there the RF phase will have changed, and you might need to retune it and/or flip the inverter on the bluebox. I know you know this but it can't hurt to mention.

Also, when you do get the CCW locked again, before you get the other PDH box form the 40m, you might want to check that you indeed have the right beam into/out of the AOM by just looking at the CW output on the TRANS CCD. If you see a 00 somewhere around an fmod of 47.5 MHz then you're golden; if you only see higher-order modes, you probably have the once-passed first order or something, like we did one time before.

Quote:

Things generally progressed yesterday up until the evening when they started un-progressing a bit.

We got the feedback to the slow channel on and working.  It uses the DAQ and has a filter with zero at 10hz and a pole at 10mHz.  We were able to see that with the cavity locked, the piezo signal was kept around 0v for longer timescale changes.

The second beam is now going through the AOM and we have maximised the power in the first order on the way through the first time, then added an aperture to pick off only the first order before reflecting the beam back through the AOM along the same line.  At the output of this we again pickoff the first order beam and use the power meter to maximise the power by changing the position of the mirror behind the AOM.  We are getting an efficiency of approx 56% from the AOM at the moment.

We then got this beam into the cavity and got it aligned well enough to see fringes and something that looked like a 00 mode flashing.  Not well aligned, but nearly there.  At this point we connected a voltage source to the modulation input on the VCO for the AOM to scan up through the cavity in the CW direction.  We will need to collect the other PDH box from the 40m.

At this point we went back to the CCW loop to lock it - the beam had somehow got misaligned.  We spent the rest of the day trying to get the beam back into alignment with the cavity without really succeeding.  We are pretty certain that none of the cavity mirrors got moved so it should still be aligned.  We only worked on steering the beam into the cavity.  We did see 00mode resonances but it doesn't want to lock on these.  We maximised the power in transmission, and were careful to check the alignment of the beam on all the PDs but to no avail.  We will look at this again today.

 

  868   Thu Jul 22 20:07:20 2010 DmassComputingDAQFB0 Problem...fixed

I ran startatf on fb0 after changing the ATF.ini file to include a channel

[controls@fb0 daq]$ startatf
audit_log_user_command(): Connection refused
atfepics: no process killed
audit_log_user_command(): Connection refused
atfepics C2 IOC Server started
./startupC2: line 2: iocC2.log: Permission denied
audit_log_user_command(): Connection refused

I did a "sudo startatf" and there were no error messages about the ./startupC2 script.

  867   Thu Jul 22 16:19:48 2010 KojiLaserPSLH1 NPRO Measured

You may have jitter because of the PZT actuation. It may cause noise in the PDH signal by some mechanism. It was difficult to measure the jitter itself in the 40m case.

Quote:

I have a question which I either don't know or can't remember the answer to:

  • Is there very little spatial distortion induced on the beam at these frequencies by using funny PZT resonances to lock, or do they exist and  we just don't care because they are out of band (presuming they aren't downconverted somehow)

 

  866   Thu Jul 22 10:13:31 2010 AlastairLaserGYROGYRO update

Things generally progressed yesterday up until the evening when they started un-progressing a bit.

We got the feedback to the slow channel on and working.  It uses the DAQ and has a filter with zero at 10hz and a pole at 10mHz.  We were able to see that with the cavity locked, the piezo signal was kept around 0v for longer timescale changes.

The second beam is now going through the AOM and we have maximised the power in the first order on the way through the first time, then added an aperture to pick off only the first order before reflecting the beam back through the AOM along the same line.  At the output of this we again pickoff the first order beam and use the power meter to maximise the power by changing the position of the mirror behind the AOM.  We are getting an efficiency of approx 56% from the AOM at the moment.

We then got this beam into the cavity and got it aligned well enough to see fringes and something that looked like a 00 mode flashing.  Not well aligned, but nearly there.  At this point we connected a voltage source to the modulation input on the VCO for the AOM to scan up through the cavity in the CW direction.  We will need to collect the other PDH box from the 40m.

At this point we went back to the CCW loop to lock it - the beam had somehow got misaligned.  We spent the rest of the day trying to get the beam back into alignment with the cavity without really succeeding.  We are pretty certain that none of the cavity mirrors got moved so it should still be aligned.  We only worked on steering the beam into the cavity.  We did see 00mode resonances but it doesn't want to lock on these.  We maximised the power in transmission, and were careful to check the alignment of the beam on all the PDs but to no avail.  We will look at this again today.

  865   Wed Jul 21 20:13:06 2010 DmassLaserPSLH1 NPRO Measured

I want to lock with the NPRO laser PZT (as done at the 40m) - and thus want the transfer functions of the NPRO PZT to phase / amplitude modulation (measured by Mott @ the 40)

  • First plot attached is PZ T to AM response

I will use the 143.8kHz dip in the AM spectrum for locking (as sidebands), then later look at the PZT error signal to find the phase transfer function. I am interested in maximizing the ratio of PM/AM. For the loop:

  • I will put in a second order passive component low pass to filter out the 2f (at 10 kHz)
  • This is quick and should allow me to take my phase transfer function and tweak the setup if needed (and consider going to that 320ish kHz dip in the AM spectrum)
  • I may be able to live with a 1 kHz UGF, as I used a 600 - 800 Hz UGF when I ran with the old frontend

 

I have a question which I either don't know or can't remember the answer to:

  • Is there very little spatial distortion induced on the beam at these frequencies by using funny PZT resonances to lock, or do they exist and  we just don't care because they are out of band (presuming they aren't downconverted somehow)
Attachment 1: H1NPRO_AM.png
H1NPRO_AM.png
  864   Wed Jul 21 20:08:53 2010 DmassComputingDAQKilled the DAQD process. It restarted automatically.

Right - I am questioning the scalability and sense of this scheme, and inquiring if this is aesthetic.... i.e. is there a reason it cannot be C2SYS for all of bridge, (possibly front end naming limitations)?

  863   Wed Jul 21 14:21:17 2010 DmassLaserPSLCharacterizing the H1 NPRO

I took some pictures of the H1(Innolight Mephisto) NPRO's scatter, and one of the neighboring lightwave NPRO for comparison

  • Iris is about 6" in front of Meph NPRO
Attachment 1: P7140293.png
P7140293.png
Attachment 2: P7140301.png
P7140301.png
Attachment 3: P7140304.png
P7140304.png
  862   Wed Jul 21 14:20:54 2010 JenneComputingDAQKilled the DAQD process. It restarted automatically.

Quote:

We are continuing the naming scheme with CX? I thought we were planning on making C2 the whole subbasement...

Like

C2ATF

C2PSL

C2TCS...

 Nope.  Each lab has its own number.  I think PSL is C3, TCS is C4, ATF gets C2....Frank has the full listing of these things.  Or at least that's the latest I heard a few months ago after several elogs back and forth.

  861   Wed Jul 21 13:59:29 2010 DmassComputingDAQKilled the DAQD process. It restarted automatically.

We are continuing the naming scheme with CX? I thought we were planning on making C2 the whole subbasement...

Like

C2ATF

C2PSL

C2TCS...

  860   Tue Jul 20 18:12:02 2010 AidanComputingDAQKilled the DAQD process. It restarted automatically.

I killed and restarted the daqd process because I wanted to add some 16Hz TCS channels to the frame builder. These are from the Athena DAQ box.

I edited the following files:

  • on fb1: /cvs/cds/caltech/chans/daq/C4TCS.ini    -the .ini file telling the frame builder what channels to record
  • on tcs_daq: /target/TCS_westbridge.db  - added the names for the ADC inputs and DAC outputs for the Athena box

 

Quote:

I added some EPICS channels to the Hartmann sensor softIoc and then added these to be recorded in the frames.

I then killed the daqd process on fb1 so it would start afresh.

  • /cvs/cds/caltech/target/softIoc/HWS.db      - the file with the HWS EPICS channels
  • /cvs/cds/caltech/chans/daq/C4TCS.ini            - the .ini file telling the frame builder what channels to record.

 

  859   Tue Jul 20 15:44:07 2010 AidanComputingGeneralFixed two new static IP addresses for TCS machines

I fixed two machines in the TCS lab to have static IP addresses on the local network.

The Athena DAQ CentOS box: 'tcs_daq' 10.0.1.34

The CentOS workstation: 'tcs_ws' 10.0.1.25

Frank, please add these to the network topology diagram you have ...

  858   Mon Jul 19 13:34:39 2010 ZachLaserGYROnoise measurement of locked cavity

I saw that you guys are planning on turning the input mirror to get a Mach-Zehnder-type thing going... any plans to do that soon? I ask because this should give us the displacement noise directly, right?

Quote:

 This is the noise of the cavity if we just lock the laser to the cavity using the fast actuator (no slow actuation) and look at the control signal to the laser.  I used a potential divider to reduce the voltage going to the spectrum analyser (I used 1.023kOhm and 9.32kOhm to give approx 10 times reduction).

The plot here is taking the V/rt(Hz) and multiplying it back up again to the full voltage, then converting to frequency using the piezo spec from the laser spec sheet of 4.11471MHz/V.

We maybe want to plot this on the main noise plot to see how far away we are from measuring a useful mirror displacement noise, so I have included the .mat file that I used to make the plot

 

  857   Sat Jul 17 01:41:20 2010 KojiLaserPSLH1 NPRO Measured

In terms of the damage of the optics, I have never seen the CVI mirror is damaged by this level of the power density.

What I have seen was a newport BK7 cube BS melting with the 10W focused on the <100um waist radius.

Quote:

This thing really didn't want its modes matched. I had to open up a can O' parameter space on it. Notas:

  • I added a static lens in front of the NPRO as well to change the mode I was working with
  • I extended the path length (will have to use some folding mirrors)
  • I checked that dwaist/dposition was dominated by one of the mirrors by about a factor of ten, so I might actually be able to fiddle with them easily once theyre in place and the PMC is locked.
  • I tried to avoid making any nasty waists on the table, the one that does exist is bigger than 100 um radius, and MAY PASS OVER WHERE I HAVE A MIRROR CURRENTLY!

 

121 um waist with 1.5 Watts is about 3*10^3 W/cm^2 - is this fine to sit pretty much on a mirror,

  • Damage thresholds for CVI coatings seem to have several to 10 "J/cm^2" for 20 nanosecond 1064 nm pulses\
    • This means it can take 500 MW/cm^2 for 20 ns - unless I know how often these pulses happen I'm not sure how to make this a CW number
    • To turn this into a normal number I assume I just do "1 J / pulse lase rate / cm^2" - I have no idea what rates are common for pulsed lasers
  • CW damage thresholds of materials seem to be order "MW/cm^2" - and if these are what's quoted by a company, it would seem that their coatings would have to not be many orders worse than this (or else the number would be meaningless) - though these companies have done other ridiculous *&!# before so...

I would appreciate any input from the peanut gallery - for now I am inclined to think that 1.5W on an HR mirror with a 240 um spot size is fine.

 

  856   Fri Jul 16 19:23:27 2010 AlastairLaserFuglynoise measurement of locked cavity

 This is the noise of the cavity if we just lock the laser to the cavity using the fast actuator (no slow actuation) and look at the control signal to the laser.  I used a potential divider to reduce the voltage going to the spectrum analyser (I used 1.023kOhm and 9.32kOhm to give approx 10 times reduction).

The plot here is taking the V/rt(Hz) and multiplying it back up again to the full voltage, then converting to frequency using the piezo spec from the laser spec sheet of 4.11471MHz/V.

We maybe want to plot this on the main noise plot to see how far away we are from measuring a useful mirror displacement noise, so I have included the .mat file that I used to make the plot

Attachment 1: freq_noise.pdf
freq_noise.pdf
Attachment 2: noise_from_locked_cavity.mat
  855   Fri Jul 16 18:54:57 2010 DmassLaserPSLH1 NPRO Measured

All the previous plots are junk - I will replace them. I made a stupid typo and propagated it through the mode matching I did, and thus nothing I did on the table made any sense.

I have mode matched to the PMC, and see what I think is TEM-00 when scanning it.

I need either an EOM, or need to use the laser PZT to put audio sidebands on the light and lock with those (as we discussed doing with the NPROs at the end stations in the VGL scheme for aLIGO / 40m)

 

EDIT - Removed old BS plot - the below plot is good!

Attachment 1: H1ModeMSoln.pdf
H1ModeMSoln.pdf
  854   Thu Jul 15 15:41:40 2010 DmassLaserPSLH1 NPRO Measured

Quote:

I unpacked the NPRO Rick sent me from the H1 PSL, and put it on the table.

I measured the total power output to be 1.574 W with the high power meter.

Waist measurements were taken with the WinCamD.

I saw what seemed to me to be a lot of junk light / scatter coming out of the NPRO. I am going ahead with putting it into the PMC in the interest of finishing.

 

The old plot was wrong. It seems I chaotically clicked blue buttons when taking my waist measurements. I attempted to mode match and noticed that what happened on the table was crazy.

[Edit - New plot added]

 I was unable to find a two lens mode matching solution using any of the available lenses in the lab constraining myself to the existing layout - I will try 3 lens solutions briefly then consider changing the optical layout. I wanted to avoid this in the interest of time, but may have to do this for the same reason.

  853   Wed Jul 14 21:48:02 2010 JennaLaserGYROCCW Lock

 Today I mounted the mode matching lenses on translation stages, and Alastair and I re-aligned the CCW beam into the cavity. We hooked up the PDH servo and brought the beam into lock. Alastair noted that it was staying locked much longer than it ever had before, but of course a slight bump to the table would bring it out of lock  (or back again). I've attached a picture of the monitor output for the locked beam.

Attachment 1: CCW_lock_714.jpg
CCW_lock_714.jpg
  852   Wed Jul 14 14:39:09 2010 DmassLaserPSLH1 NPRO Measured

I unpacked the NPRO Rick sent me from the H1 PSL, and put it on the table.

I measured the total power output to be 1.574 W with the high power meter.

Waist measurements were taken with the WinCamD.

I saw what seemed to me to be a lot of junk light / scatter coming out of the NPRO. I am going ahead with putting it into the PMC in the interest of finishing.

 

The old plot was wrong. It seems I chaotically clicked blue buttons when taking my waist measurements. I attempted to mode match and noticed that what happened on the table was crazy.

[Edit - New plot added]

Attachment 1: NPROWaist.pdf
NPROWaist.pdf
  851   Mon Jul 12 18:48:15 2010 JennaMiscGYROTranslation Stages from 40m

 These translation stages are from a cabinet on the x-arm of the 40m. Rana is ordering some new stages from Newport, but in the meantime I'll use these to optimize the lens positions.

Attachment 1: HPIM3679.JPG
HPIM3679.JPG
  850   Mon Jul 12 13:16:59 2010 JennaLaserGYROCCW Alignment

 After I installed the mode matching lenses last week, I aligned the CCW beam into the cavity. The solid red lines in the picture below are the incoming beam, and I've included a picture of the CCD camera output on the monitor and the photodetector output on the oscilloscope. The beam is currently being scanned at .1 Hz. I've also aligned the beam coming out (dashed red line) and through the Faraday isolator and into the photodetector. I haven't had a chance yet to look at the output of this though.

Attachment 1: 7-9_CCW_alignment.png
7-9_CCW_alignment.png
Attachment 2: 7-9_CCW_alignment-2.png
7-9_CCW_alignment-2.png
  849   Thu Jul 8 13:17:01 2010 DmassLaserDoublingForty Two

Alternately, I can write the really simple equation for phase noise...

equation 5.24a from the Bloembergen paper (non depleted pump approx with plane waves):

Phi2 = 2*Phi1-pi/2+deltak*z/2

and just write the phase noise in terms of the phase mismatch. Note that this can be in terms of temperature.

  848   Thu Jul 8 00:16:43 2010 DmassLaserDoublingForty Two

I think I have found the answer to the doubling phase noise after a few pages of algebra and differential equations. I will have to confirm this at a later date, but for now I assume it's correct:

When I lock the phase of the Mach Zehnder in the IR, the phase of the green is:

Phi = Constant + w1/c*[(n1-n2)*dL/dT+(dn1/dT-dn2/dT)*L]*delta_T

  • w1 is the frequency of the field @ 1064 nm (radians / sec)
  • L is the length of the crystal
  • n1 is the refractive index of PPKTP @ 1064 nm
  • n2 is the refractive index of PPKTP @ 532 nm
  • delta_T is the temperature fluctuation at the crystal

In summary for the "locked Mach Zehnder scheme"

If I treat the phase of the IR as my error signal, and push it to zero (or to some constant), the fluctuations in my green (and thus sensing noise), will be dominated by temperature fluctuations of the oven.

  • If this is true, any suppression of temperature noise should be seen as a reduction in the phase noise of the green when the Mach Zehnder is locked.
  847   Wed Jul 7 13:59:26 2010 JennaLaserGYROMode Matching

 The lenses I'm using for the mode matching telescope are Newport KPX103 AR.18 (175mm) and KPX097 AR.18 (125mm). I contacted to the company to check how the focal length would change for 1064nm. These focal lengths are calculated for a wavelength of 589nm, but the index of refraction for the BK7 glass they're made of decreases from 1.51673 at 589nm to 1.05663 at 1064nm, so the focal lengths increase by 2% (assuming a thin lens).

If I left the lenses in their current positions which were calculated for 125 and 175mm, this gives a power coupling value of eta=.7674, so I have calculated the new distances for the adjusted focal lengths of 127.5mm and 178.5m. The 125mm lens should now be located 597.42mm from the mirror at the corner of the table (4w2l), and the 175mm lens should be 345.21mm from that. These positions give eta=.9910.

I've plotted how eta changes as a function of the lens positions as well. Lens 1 is the 125mm (or rather 127.5mm lens) and Lens 2 is the 175/178.5m. These plots cover +/- 10mm from the calculated ideal position. For lens 1 this value is 597.42mm, as quoted earlier, and for lens 2 it is 2383.46mm, which is the distance from lens 2 to the input mirror.

Attachment 1: power_coupling_vs_lens_2_position_for_1064nm.png
power_coupling_vs_lens_2_position_for_1064nm.png
Attachment 2: power_coupling_vs_lens_1_position_for_1064nm.png
power_coupling_vs_lens_1_position_for_1064nm.png
  846   Tue Jul 6 21:07:42 2010 ranaThings to BuyGYROIR viewer missing: found in TCS lab

We wasted some time today looking around for the IR viewer. We found that it had been taken into the TCS lab with no notice left about it being borrowed.

This is unacceptable and the second time this has happened. The next time anyone from the TCS lab tries to borrow equipment, refuse.

Tell whomever it is to come and talk to me if they have a problem with this policy.

  845   Thu Jul 1 17:47:19 2010 JennaLaserGYROMode Matching Solution

Here is the spherical lens solution that doesn't take into account the astigmatism of the cavity. All of these values were calculated using the programs that I posted in my earlier mode matching post.

The required beam waist size for the cavity at the input mirror is 931µm. I got this number by calculating the X and Y components of the waist separately and then averaging them. 

The mode matching solution I get from the program is to use a 125mm and a 175mm lens. The 125mm lens should be located 232.292mm from the mirror in the corner of the table (at 4W2L), and the 175mm lens 337.616mm from that.  This lens solution gives a beam waist size of exactly 931µm, and an offset from the input mirror of 1.7µm. Taking into account the astigmatism of the cavity, this gives an overlap function value of η=.9910.

This eta value is still very high even with the astigmatism of the cavity, and doesn't change much until you move the lenses dramatically. I need to modify one of my programs to plot the eta value as the lens position changes, and I'll add a plot once I get that written to better quantify this. 

 

 

  844   Wed Jun 30 18:19:47 2010 JennaLaserGYROCavity Mirrors

 Alastair and I went down to the lab today and changed the input mirror in the cavity to a two inch mirror and switched out the 6m mirrors with 9m ones. The input mirror has a transmission of 6515ppm, and the output mirror has a transmission of 5700ppm.

  843   Wed Jun 30 00:35:28 2010 ranaLaserGYROComparison of frequency noise and mechanical noise for the gyro cavity

I think its still worth measuring. If you really measure just the frequency noise of the laser that's great. For the second measurement you just rotate the output mirror by 90 deg and make a balanced homodyne Mach Zender measurement of the mirror motions which rejects frequency noise.

  842   Tue Jun 29 17:48:31 2010 ZachLaserGYROComparison of frequency noise and mechanical noise for the gyro cavity

Is there any reason to assume that it will be quieter on our table than it is on the 40m PSL table? I would imagine that our top-notch temperature control system ruins us at low frequency, while the chiller and all our other equipment must do wonders for us in the audio band.

Quote:

It looks like the laser frequency noise will be slightly too high to measure the vibration of the mirrors in the cavity using the control signal for the locking, especially if our cavity is quieter than the MZ at the 40m (which we are hoping it will be).  It would still be nice to get a measurement of the mechanical noise of the mirrors somehow.

 

Quote:

 Here is a plot of laser frequency noise from a Lightwave NPRO 126 laser converted into m*Hz-1/2 for comparison with mechanical noise taken from the MZ at the 40m.

 

 

  841   Tue Jun 29 15:41:01 2010 AlastairLaserGYROComparison of frequency noise and mechanical noise for the gyro cavity

It looks like the laser frequency noise will be slightly too high to measure the vibration of the mirrors in the cavity using the control signal for the locking, especially if our cavity is quieter than the MZ at the 40m (which we are hoping it will be).  It would still be nice to get a measurement of the mechanical noise of the mirrors somehow.

 

Quote:

 Here is a plot of laser frequency noise from a Lightwave NPRO 126 laser converted into m*Hz-1/2 for comparison with mechanical noise taken from the MZ at the 40m.

 

  840   Tue Jun 29 14:27:10 2010 JennaLaserGYROComparison of frequency noise and mechanical noise for the gyro cavity

 Here is a plot of laser frequency noise from a Lightwave NPRO 126 laser converted into m*Hz-1/2 for comparison with mechanical noise taken from the MZ at the 40m.

Attachment 1: frequency_and_mechanical_noise.png
frequency_and_mechanical_noise.png
  839   Tue Jun 29 13:12:56 2010 AlastairLaserGYROMode Matching

Quote:

When posting MATLAB code to the elog - attach as an archive so it doesn't take up many pages when viewing all posts in the elog

 done

ELOG V3.1.3-